OpenStudy (anonymous):

write and equation for the following transformation of y=x' a vertical stretch by a factor of 4

OpenStudy (anonymous):

it would be y-4=x'

OpenStudy (anonymous):

how did you get that though? its a workpad. im confused.

OpenStudy (anonymous):

okay let me clarify . have you got question ?

OpenStudy (anonymous):

i just am not sure how you got that. if its a vertical stretch does that mean you just put a subtraction sign?

OpenStudy (anonymous):

nope i got my interpretation of question its like this you got graph of y=x' and now you wanna stretch that graph in vertical direction by 4 units here we have a basic rule if you wanna shift the graph of y=f(x) by 'a' units what you do is subtract a from y so your equation of new shifted graph is y-a=f(x) now you know why i subtracted 4 from y

OpenStudy (anonymous):

oh ok that makes more since. thank you for your help :)

OpenStudy (anonymous):

you are most welcome !! :)